# WaterPark: A Robustness Assessment of Language Model Watermarking

### v2

@@ Abstract
WaterPark: A Robustness Assessment of Language Model Watermarking
@@ 1 Introduction
Previous Conclusion Refined Conclusion Explanation Consistency
@@
We develop WaterPark, the first open-source platform dedicated to evaluating the attack robustness of LLM watermarkers in a unified and comprehensive manner. As of 12/15/2024, WaterPark integrates 10 state-of-the-art watermarkers, 12 representative watermark removal attacks, and 8 key metrics. Moreover, WaterPark offers a comprehensive suite of tools for in-depth robustness assessment, including next-token distribution comparisons, attack combination analyses, and what-if scenario evaluations. Leveraging WaterPark, we empirically evaluate the attack resilience of representative LLM watermarkers, leading to many interesting findings, which challenge the conclusions in prior work, as summarized in Table 1. We also explore how a watermarker’s design choices impact its attack robustness, unveiling critical trade-offs between different types of robustness.
@@ 2 LLM Watermarking
Typically, the generator produces watermarked texts iteratively. Note that in symmetric schemes, $k=k^{\prime}$, while in asymmetric schemes (Fairoze et al., 2023), $k\neq k^{\prime}$. This study mainly focuses on symmetric watermarking schemes due to their widespread adoption.
@@ 4.1 Experimental Setting
Our evaluation considers all the watermarkers in Table 6 and attacks in Table 2. We employ two LLMs OPT-1.3B and Llama2-7B-chat-hf to represent small and large models respectively. We use two datasets C4 and HC3 to simulate diverse tasks such as question-answering and text completion.
We implement and configure all watermarkers and attacks in accordance with their official documentation. Table 5 summarizes the default parameter settings. To validate our implementation, we assess their effectiveness and fidelity, as reported in §E and §F, corroborating results from existing literature. A detailed description of each attack can be found in §D. We measure the TPRs of different watermarkers against each attack (with their FPRs fixed as 1%).
Attack Category Resource
Lowercasing Linguistic variation /
Contracting /
Expanding /
Misspelling Lexical editing Common misspellings
Typoing /
Synonymizing WordNet
Swapping /
Copy-pasting Text-mixing Non-watermarked text
Deep-paraphrasing Paraphrasing LLM-based paraphraser
Translating LLM-based translator
Black-box adversarial attack Generic detector
Table 2: A taxonomy of watermark removal attacks.
4.2 Robustness – Observational Study
We systematically assess the resilience of existing LLM watermarkers against representative watermark removal attacks. All the results are shown in Table  3.
@@ 4.2 Robustness – Observational Study
UB 1.000 1.000 1.000 0.946 0.030 0.088 0.992 0.996 0.002 0.009 0.048 0.410 0.461 0.149 0.306
@@ 4.2.1 Linguistic Variation Attack
i) Overall, most watermarkers show strong resilience against linguistic variation attacks. For instance, TGRL reaches close to 100% TPRs under all three attacks. ii) UPV is the only watermarker marginally susceptible to such attacks. This can be explained by the fact that its neural network-based detector primarily depends on implicit textual features, which appear to be sensitive to changes in linguistic characteristics.
@@ 4.2.2 Lexical Editing Attack
i) TRGL, UB, SIR and GO tend to be more vulnerable to lexical editing attacks, compared with RDF and UG watermarkers. Intuitively, as text-dependent watermarkers (e.g., TRGL, UB, and GO) use previous tokens as the context for the next token in both the watermark generator and detector, the lexical editing thus causes a mismatch between the generator and detector. ii) UB exhibits much higher vulnerability to such attacks, compared with the others. For example, its TPR drops near zero under the typoing and swapping attacks. This may be explained as follows. Recall that, to achieve unbiasedness, UB applies “hard” perturbation on the next-token distribution (e.g., by rejecting half of the vocabulary); thus the disruption to the previous tokens tends to cause a more significant mismatch between the generator and detector, compared with other watermarkers that employ “soft” perturbation (e.g., distribution shift and transform). iii) The SIR shows notably inferior performance, especially when subjected to typoing attacks. While it employs a neural network to predict token-specific logit perturbations that are designed to be unbiased (no preference for particular tokens) and balanced (with perturbations summing to zero), its results are suboptimal, suggesting that it is fundamentally challenging to make accurate predictions under such perturbations. iv) The typoing attack is particularly effective, as it may cause significant tokenization errors.
@@ 4.3 Robustness – Causal Analysis
In addition to the observational studies, we further consider conducting causal analysis to understand the impact of individual design choices (e.g., samplers). However, this is challenging in our context because the various components of a watermarker are often highly interconnected and difficult to decouple. For example, the distribution-reweight strategy and the difference-based detection of UB are closely linked and cannot be easily replaced by other designs. To address this challenge, we select two watermarkers with their only difference in the design of one specific component (e.g., context dependency). The results are shown in Figure 3.
Figure 3: Watermarker robustness to multi-attacks. a) Context dependency: TGRL (text-dependent) and UG (context-free); b) Generation strategy: TGRL (distribution-shift) and GO (distribution-transform); c) Detection method: UPV (Model-based) and UPV-key (Score-based).
@@ 4.3.1 Context Dependency
We select TGRL and UG to represent text-dependent and context-free designs respectively. TGRL uses the previous $k$ tokens as a randomness seed to divide the vocabulary into red/green lists for the current token, whereas UG uses fixed red/green lists in generating all the tokens. In general, UG performs similarly to TGRL. However, when subjected to the Dipper attack at varying intensity levels, UG consistently outperforms TGRL. For instance, as we increase the attack intensity (e.g. DP-60-20, DP-60-40), which involves rearranging the text, UG maintains a relatively stable TPR with only minor reductions. In contrast, TGRL’s TPR significantly drops, particularly at DP-60-20. Recall that the Dipper attack extensively paraphrases and reorders the text, leading to substantial changes in consecutive tokens. Due to its text-dependent design, TGRL struggles to maintain consistency in random seeds between detection and generation. In contrast, UG’s context-free design avoids this issue. This aligns with the results in UG (Zhao et al., 2023), showing that when subjected to paraphrasing attacks, UG consistently outperforms TGRL.
@@ 4.3.2 Generation Strategy
TGRL and GO are both context-dependent watermarkers. As it relies on distribution shift generation, TGRL tends to be less robust against copy-paste attacks. In contrast, although copy-paste attacks can significantly disrupt detection and dilute the watermark by inserting large amounts of text, due to its distribution transform design, GO can better maintain the watermark’s concentration. As a result, GO achieves a higher detection rate of watermarked text after such attacks.
@@ 4.3.3 Detection Method
UPV employs model-based detection to detect watermarks. Alternatively, it can also use score-based detection similar to TGRL. The original paper claims that model-based detection is more resistant to paraphrasing attacks than score-based detection. However, our results indicate that this advantage is not significant. Furthermore, score-based UPV demonstrates stronger robustness against weaker attacks (e.g., misspelling and swapping attacks). Using model-based detection introduces higher uncertainty compared to score-based detection and does not provide watermarkers with greater flexibility in resisting paraphrasing attacks.
Both RDF and GO utilize distribution-transform generation and rely on score-based detection, with RDF specifically employing edit score. RDF shows marginally higher robustness against lexical editing attacks with fixed token length compared to GO. However, it exhibits lower resilience against high-intensity text-mixing attacks (e.g., CP-3-10), where token length varies considerably.
@@ 5 Discussion
For each watermarker, we mainly use its specific detector to detect watermarked texts. Here, we explore a generic, neural network-based detector as an alternative. To this end, we employ a pre-trained RoBERTa model and fine-tune it as a binary classifier using watermarked and non-watermarked texts. We use OPT-1.3B as the underlying LLM and C4 as the reference dataset. The watermarked text is generated by the LLM and the watermarker jointly, whereas the non-watermarked text is produced by ChatGPT-3.5-Turbo using the same prompt to mimic human response. Table 4 shows the TPRs of watermarkers with generic detectors (with FPRs fixed as 1%).
@@ 5.2.1 Varying Attack Intensity
Figure 5: Resilience of watermarkers against increasingly intensive Dipper attacks.
One straightforward way to improve an attack’s effectiveness is to increase its intensity, potentially at the cost of other metrics (e.g., text quality). Here, we consider the Dipper attack (Krishna et al., 2023) under varying intensity settings, denoted as DP-$l$-$o$, where lexical change ($l$) indicates that $l$of the given text is paraphrased and order change ($o$) indicates that $o$of the text is re-ordered. We compare the watermarkers’ resilience under varying attack intensity, as shown in Figure 5.
@@
Figure 6: Resilience of watermarkers against individual (left) and combined (right) attacks.
@@ 5.2.2 Combining Simple Attacks
We first explore whether combining two attacks improves the attack’s effectiveness. Here, considering the most feasible combinations, we only combine two simple attacks from the “weak” linguistic variation and lexical editing attacks. Figure 6 illustrates the effectiveness of such combined attacks.
@@ 5.2.3 Adaptive Attack
Figure 9: Effectiveness of paraphrasing attacks with ChatGPT-3.5-Turbo as the paraphraser.
@@ 5.2.4 Leveraging Expert LLMs
[…]
We have the following findings. i) The detection rates of all the watermarkers drop below 0.3 after one round of GPT paraphrasing, indicating that using highly capable LLMs to paraphrase watermarked texts is a dominant attack that effectively nullifies most watermarkers. However, as specified in our threat model (§3.1), access to highly capable LLMs may fall outside the scope of our robustness assessment, as their availability negates the need for watermark removal attacks. ii) For watermarkers that survive the first round of paraphrasing (e.g., UG and GO), their detection rates quickly drop below 0.15 as the adversary applies multiple rounds of paraphrasing. This observation corroborates the findings in (Zhang et al., 2023).
@@ Appendix D Details of Watermark Removal Attacks
Figure 10: Overall effectiveness of different watermarkers in generating and detecting watermarked texts: ROC of (a) OPT-C4, (b) OPT-HC3, (c) Llama2-C4, and (d) Llama2-HC3; TPR (with FPR fixed as 0.01) on (e) C4 and (f) HC3.
@@ E.1 Overall Effectiveness
Meanwhile, most methods observe marginal performance drops on Llama2-7B, as shown in Figure 10 (c-d). For instance, compared with its performance on OPT-1.3B, the AUC of SIR drops by 0.11 and 0.08 on C4 and HC3 respectively. This observation aligns with previous research  (Kuditipudi et al., 2023), indicating that watermarkers are more effective on OPT compared with Llama2. This phenomenon can be partly understood through the following explanation. In contrast of less capable LLMs (e.g., OPT-1.3B), Llama2-7B typically produces texts of lower perplexity. Since most watermarkers inject watermarks by slightly altering the next-token distribution, the lower perplexity in Llama2’s outputs hampers the effectiveness of such perturbations. Moreover, it is observed that the performance of various methods varies significantly across different datasets. For instance, UG’s AUC differs by 0.28 between C4 and HC3, while UB’s AUC differs by 0.11. This observation is further supported by the TPR measures at fixed FPRs (fixed as 1%), as shown in Figure 10 (e-f). Our findings suggest that the concrete dataset/task tends to have a larger impact on watermarkers over more capable LLMs.

### v3

# Jiacheng Liang

@@ Abstract
Watermark under Fire: A Robustness Evaluation of LLM Watermarking
@@ 1 Introduction
Previous Conclusion Refined Conclusion Explanation Consist.
@@
We develop WaterPark, the first open-source platform dedicated to evaluating the attack robustness of LLM watermarkers in a unified and comprehensive manner. As of 5/15/2025, WaterPark integrates 10 state-of-the-art watermarkers, 12 representative watermark removal attacks, and 8 key metrics. Moreover, WaterPark offers a comprehensive suite of tools for in-depth robustness assessment, including next-token distribution comparisons, attack combination analyses, and what-if scenario evaluations. Leveraging WaterPark, we empirically evaluate the attack resilience of representative LLM watermarkers, leading to many interesting findings, which challenge the conclusions in prior work, as summarized in Table 1. We also explore how a watermarker’s design choices impact its attack robustness, unveiling critical trade-offs between different types of robustness. In addition, we provide detailed deployment and evaluation guidelines in  J and  I.
@@ 2 LLM Watermarking
Typically, the generator produces watermarked texts iteratively. At each iteration, with access to a secret key $k$, the previous tokens $x_{<t}$, and the LLM’s next-token distribution $p(x_{t}|x_{<t})$, the generator generates a perturbed distribution $\tilde{p}(x_{t}|x_{<t})$, from which the next token $x_{t}$ is sampled. Meanwhile, with access to a secret key $k$, the detector determines whether $M$ generates a given text.
@@ 4.1 Experimental Setting
To comprehensively evaluate the robustness of LLM watermarkers, we assess all methods listed in Table  7 against the full range of attacks defined in Table  8.
We include three LLMs of varying scales(i.e. OPT-1.3B (small), LLaMA3-7B-chat (medium), and Qwen2.5-14B-Instruct (large)) and five datasets spanning different styles and domains: C4 (text completion), HC3 (QA), Story Completion (creative writing), Law Stack Exchange (legal), and Paper Conclusion (academic summarization). This design ensures broad and unbiased robustness assessment.
All watermarkers and attacks are implemented following their official documentation. Table 6 summarizes the default parameter configurations used in our experiments. We further validate the fidelity and effectiveness of our implementations in Sections  E and  F, confirming consistency with results reported in the original works. The evaluation results across various datasets are presented in Table 10. The influence of generation length is discussed in Section  E.2, and detailed descriptions of each attack type are provided in Appendix  D.
We report the true positive rates (TPRs) of all watermarking methods under each attack type, with the false positive rate (FPR) consistently fixed at 1%.
@@
TGRL 0.993 0.944 0.944 0.831 0.429 0.222 0.887 0.915 0.000 0.667 0.833 0.833 0.667 0.485 0.222
UG 0.993 0.857 0.833 0.976 0.929 0.930 0.976 0.738 0.857 0.714 0.714 0.991 0.976 0.877 0.921
UPV 0.400 0.400 0.400 0.430 0.080 0.000 0.380 0.360 0.000 0.000 0.000 0.000 0.200 0.080 0.240
RDF 0.999 0.998 0.996 0.996 0.976 0.979 0.991 0.993 0.872 0.893 0.932 0.978 0.905 0.738 0.978
UB 0.980 1.000 1.000 0.962 0.000 0.033 0.921 1.000 0.018 0.042 0.000 0.485 0.514 0.103 0.255
SIR 0.978 0.580 0.500 0.420 0.320 0.320 0.440 0.560 0.340 0.360 0.360 0.460 0.400 0.300 0.000
GO 0.996 0.996 0.996 0.994 0.982 0.956 0.986 0.996 0.864 0.887 0.946 0.982 0.640 0.560 0.667
4.2 Robustness – Observational Study
Our comprehensive cross-model and cross-dataset analysis, as detailed in Table  9, demonstrates a consistent ranking of robustness across various models and datasets. This consistency indicates that the inherent algorithmic design of the watermarkers is the primary factor influencing their relative robustness, rather than model-specific or dataset-specific factors.
To provide a thorough evaluation, we selected the Qwen2.5-14B-Instruct model with the Paper Conclusion Dataset as our primary evaluation setting. As shown in Table  2, this combination effectively captures the observed robustness patterns, making it a reliable basis for our main results. This choice reflects the broader trends identified in our extensive cross-analysis efforts.
@@ 4.2.1 Linguistic Variation Attack
i) Overall, most watermarkers show strong resilience against linguistic variation attacks. For instance, TGRL reaches close to 100% TPRs under all three attacks. ii) UPV is the watermarker marginally susceptible to such attacks. This can be explained by the fact that its neural network-based detector primarily depends on implicit textual features, which appear to be sensitive to changes in linguistic characteristics. iii) The SIR shows notably inferior performance. While it employs a neural network to predict token-specific logit perturbations that are designed to be unbiased (no preference for particular tokens) and balanced (with perturbations summing to zero), its results are suboptimal, suggesting that it is fundamentally challenging to make accurate predictions under such perturbations.
@@ 4.2.2 Lexical Editing Attack
i) TRGL, UB, SIR, and UPV tend to be more vulnerable to lexical editing attacks, compared with RDF and UG watermarkers. Intuitively, as text-dependent watermarkers (e.g., TRGL, UB, SIR, and UPV) use previous tokens as the context for the next token in both the watermark generator and detector, the lexical editing thus causes a mismatch between the generator and detector. ii) UB exhibits much higher vulnerability to such attacks, compared with the others. For example, its TPR drops near zero under the typoing and swapping attacks. This may be explained as follows. Recall that, to achieve unbiasedness, UB applies “hard” perturbation on the next-token distribution (e.g., by rejecting half of the vocabulary); thus the disruption to the previous tokens tends to cause a more significant mismatch between the generator and detector, compared with other watermarkers that employ “soft” perturbation (e.g., distribution shift and transform). iii) UPV and SIR exhibit similar vulnerabilities to the previous attack
@@ 4.3 Robustness – Causal Analysis
In addition to the observational studies, we further consider conducting causal analysis to understand the impact of individual design choices (e.g., samplers). However, this is challenging in our context because the various components of a watermarker are often highly interconnected and difficult to decouple. To address this challenge, we select two watermarkers with their only difference in the design of one specific component (e.g., context dependency). The results are shown in Figure 3.
Figure 3: Watermarker robustness to multi-attacks. a) Context dependency: TGRL (text-dependent) and UG (context-free); b) Generation strategy: TGRL (distribution-shift) and GO (distribution-transform); c) Detection method: UPV (Model-based) and UPVstat (Score-based). RDF and GO.
@@ 4.3.1 Context Dependency
We select TGRL and UG to represent text-dependent and context-free designs, respectively. TGRL uses the previous $k$ tokens as a randomness seed to divide the vocabulary into red/green lists for the current token, whereas UG uses fixed red/green lists across all positions.
As shown in (a), UG consistently outperforms TGRL under most attack categories. This includes not only strong paraphrasing attacks such as DP-60 and DP-40, but also lexical editing attacks (e.g., Typo, Syno) and text-mixing attacks (e.g., CP1-10, CP3-10). In contrast, TGRL exhibits a noticeable drop in true positive rate (TPR) under these challenging perturbations, especially when token ordering and surface forms are significantly altered.
This performance discrepancy is expected. Attacks like Dipper-style paraphrasing and text mixing disrupt the continuity and order of input tokens, breaking the alignment assumed by TGRL’s context-dependent randomness during detection. In contrast, UG’s context-free design applies consistent token-level watermarking that remains robust regardless of token rearrangement or local corruption. These results further validate prior observations in UGZhao et al. (2023) regarding its robustness to syntax-level transformations.
@@ 4.3.2 Generation Strategy
Both TGRL and GO adopt context-dependent watermarking. However, they differ in their generation mechanisms: TGRL employs a distribution-shift approach with soft constraints, whereas GO utilizes a deterministic distribution-transform method on token logits.
As observed in the (b), GO consistently outperforms TGRL under text-mixing attacks, including CP1-10, CP3-10, and CP3-25. While TGRL’s true positive rate (TPR) drops significantly under these perturbations, GO maintains high detection performance across all copy-paste scenarios. GO also shows moderate robustness under paraphrasing attacks such as DP-40 and DP-60, where TGRL again underperforms.
The key difference lies in signal strength: GO’s transform-based sampling embeds consistent token-level signals resilient to inserted distractors. TGRL’s soft sampling is more vulnerable to dilution and misalignment from non-watermarked fragments.
@@ 4.3.3 Detection Method
UPV supports both model-based and score-based detection. While the original work suggests model-based detection improves robustness against paraphrasing, our results challenge this claim. As shown in the radar plot (right), both UPV (model-based) and UPVstat (score-based) exhibit near-zero TPR under strong paraphrasing (DP-60, DP-40) and text-mixing (CP3-25), indicating a critical failure in handling high-intensity attacks. Interestingly, UPVstat slightly outperforms UPV under milder perturbations such as misspelling and synonym substitution. This implies that model-based detection introduces additional variance and is more prone to subtle distributional shifts, whereas score-based detection retains better stability in cleaner settings. However, neither variant withstands red-team scenarios effectively, calling into question the overall practicality of UPV’s detection mechanism.
Both RDF and GO adopt distribution-transform generation with score-based detection, but differ in scoring criteria. RDF leverages edit-based scoring and demonstrates consistently higher robustness than GO across most attacks. In particular, RDF shows greater resilience under strong paraphrasing (DP-40, DP-60) and text rearrangement (CP3-25), highlighting the benefits of edit-distance awareness in detecting structurally altered text.
@@ 5 Discussion
For each watermarker, we mainly use its specific detector to detect watermarked texts. Here, we explore a generic, neural network-based detector as an alternative. To this end, we employ a pre-trained RoBERTa model and fine-tune it as a binary classifier using watermarked and non-watermarked texts. We use OPT-1.3B as the underlying LLM and C4 as the reference dataset. The watermarked text is generated by the LLM and the watermarker jointly, whereas the non-watermarked text is produced by ChatGPT using the same prompt to mimic human response. Table 3 shows the TPRs of watermarkers with generic detectors (with FPRs fixed as 1%).
@@ 5.2.1 Varying Attack Intensity
Table 4: Robustness of Watermarkers against Dipper Attacks with Varying Lexical and Order Changes.
Water marker Attack Intensity(Dipper - lexical ($l$) - order ($o$) )
20 40 60 40-20 60-20 60-40
TGRL 0.952 0.858 0.550 0.746 0.406 0.546
UG 0.940 0.916 0.848 0.894 0.846 0.789
UPV 0.296 0.057 0.645 0.742 0.630 0.606
RDF 0.988 0.962 0.840 0.914 0.686 0.714
UB 0.461 0.149 0.025 0.070 0.016 0.009
SIR 0.902 0.814 0.050 0.065 0.067 0.051
GO 0.971 0.847 0.500 0.785 0.378 0.460
One straightforward way to improve an attack’s effectiveness is to increase its intensity, potentially at the cost of other metrics (e.g., text quality). Here, we consider the Dipper attack (Krishna et al., 2023) under varying intensity settings, denoted as DP-$l$-$o$, where lexical change ($l$) indicates that $l$of the given text is paraphrased and order change ($o$) indicates that $o$of the text is re-ordered. We compare the watermarkers’ resilience under varying attack intensity, as shown in Table 4.
@@ 5.2.2 Combining Simple Attacks
Table 5: Resilience of watermarkers against individual and combined attacks(compared to single attack).
Water marker Contra + Swap + Syno +
Typo Swap Low Low Missp Typo Swap Low Typo
TGRL -.046 -.008 -.002 -.232 -.122 -.648 -.086 -.020 -.268
UG -.034 .006 -.009 -.066 -.002 -.080 -.008 -.105 -.125
UPV -.192 -.437 -.415 -.505 -.497 -.505 -.548 -.534 -.552
RDF -.021 .043 .009 .013 .039 -.267 -.004 -.002 -.132
UB -.028 -.016 -.022 -.028 -.026 -.088 -.982 -.258 -.956
SIR -.524 -.858 -.840 -.838 -.866 -.890 -.862 -.856 -.894
GO -.068 .006 .000 -.258 -.092 -.834 -.098 -.002 -.118
We first explore whether combining two attacks improves the attack’s effectiveness. Here, considering the most feasible combinations, we only combine two simple attacks from the “weak” linguistic variation and lexical editing attacks. Table 5 illustrates the effectiveness of such combined attacks.
@@ 5.2.4 Leveraging Expert LLMs
We investigate the impact of highly capable LLMs in paraphrasing attacks by using ChatGPT to rewrite watermarked texts. For each watermarker, we randomly sample 100 successfully detected watermarked texts and query ChatGPT with prompt: “Paraphrase the following text and keep length similar”. To assess the effect of paraphrasing strength, we also apply up to 5 rounds of iterative paraphrasing, following the strategy in (Zhang et al., 2023), which originally tested up to 300 rounds using T5. We focus on practical, resource-limited attackers, while their work (Zhang et al., 2023) explores worst-case robustness. Results (Figure  7) show that: i) A single round of ChatGPT paraphrasing reduces detection rates of all watermarkers below 0.3, demonstrating that expert LLMs pose a strong threat. However, as noted in §3.1, such powerful LLMs may fall outside our target threat model. ii) For more resilient methods (e.g., UG and GO), detection rates still drop below 0.15 after a few paraphrasing rounds, confirming similar trends observed in (Zhang et al., 2023).
Figure 7: Effectiveness of paraphrasing attacks with ChatGPT as the paraphraser.
@@ Appendix D Details of Watermark Removal Attacks
Attack Category Resource
Lowercasing Linguistic variation /
Contracting /
Expanding /
Misspelling Lexical editing Common misspellings
Typoing /
Synonymizing WordNet
Swapping /
Copy-pasting Text-mixing Non-watermarked text
Deep-paraphrasing Paraphrasing LLM-based paraphraser
Translating LLM-based translator
Black-box adversarial attack Generic detector
Table 8: A taxonomy of watermark removal attacks.
@@
Figure 8: Overall effectiveness of different watermarkers in generating and detecting watermarked texts: ROC of (a) OPT-C4, (b) OPT-HC3, (c) Llama3-C4, and (d) Llama3-HC3; TPR (with FPR fixed as 0.01) on (e) C4 and (f) HC3.
@@ E.1 Overall Effectiveness
Meanwhile, most methods observe marginal performance drops on Llama3-7B, as shown in Figure 8 (c-d). For instance, compared with its performance on OPT-1.3B, the AUC of SIR drops by 0.11 and 0.08 on C4 and HC3 respectively. This observation aligns with previous research  (Kuditipudi et al., 2023), indicating that watermarkers are more effective on OPT compared with Llama3. This phenomenon can be partly understood through the following explanation. In contrast of less capable LLMs (e.g., OPT-1.3B), Llama3-7B typically produces texts of lower perplexity. Since most watermarkers inject watermarks by slightly altering the next-token distribution, the lower perplexity in Llama3’s outputs hampers the effectiveness of such perturbations. Moreover, it is observed that the performance of various methods varies significantly across different datasets. For instance, UG’s AUC differs by 0.28 between C4 and HC3, while UB’s AUC differs by 0.11. This observation is further supported by the TPR measures at fixed FPRs (fixed as 1%), as shown in Figure 8 (e-f). Our findings suggest that the concrete dataset/task tends to have a larger impact on watermarkers over more capable LLMs.